# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v3

@@ Acknowledgments
This bound is true for any fixed unit $\mathbf{r}$. In order to extend the bound simultaneously for all unit $\mathbf{r}$, we will use a standard $\epsilon$-net argument: Fix some $\epsilon>0$, and let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net of the unit Euclidean ball in $\mathbb{R}^{T}$, of size $\leq\left(\frac{2}{\epsilon}+1\right)^{T}$ (namely, a collection of unit vectors so that each $\mathbf{r}$ in the unit ball is $\epsilon$-close in Euclidean distance to one of the vectors). The existence of an $\epsilon$-net of this size is shown, for example, in [Vershynin, 2012, Lemma 5.2]. Using a union bound and the displayed equation above, it follows that with probability at least $1-\left(\frac{2}{\epsilon}+1\right)^{T}\exp\left(-\frac{n}{30}\right)$,
@@
Combining Lemma A.1 and Lemma A.2 with a union bound, and applying them to Eq. (6), we get the following: With probability at least $1-\delta-\exp\left(T\log(2n)-\frac{n}{32}\right)$ over $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{T}$, we have that for any $\mathbf{w}$, there is some subset $\mathcal{L}\subseteq[n]$ of size $|\mathcal{L}|\geq\frac{n}{20}$ for which the following inequalities hold:
@@
where $(1)$ is by Lemma A.1, $(2)$ is by the fact that a minimum can be upper bounded by an average, $(3)$ is by Lemma A.2 and the fact that $|\mathcal{L}|\geq\frac{n}{20}$, $(4)$ is by the assumption in the proposition statement, and $(5)$ is by definition of the vector $\mathbf{r}$ and scalar $r_{0}$ in Eq. (5).
@@
A.5 Proof of Theorem 4.2
Examining the conditions in Proposition 4.6, Lemma 4.4 and Lemma 4.5, we see that in order for all of them to be applicable, we must choose $\alpha\lesssim\frac{\beta}{\sqrt{T}}$ (Proposition 4.6), $\beta\lesssim\sqrt{\frac{1}{T\log(n)}}$ (Lemma 4.5), yet the margin guarantee is $\approx\frac{\alpha}{\sqrt{T}}$ (Lemma 4.4). Thus, to make the margin $\frac{\alpha}{\sqrt{T}}$ as large as possible, yet satisfying all of the constraints, we should choose $\beta\approx\sqrt{\frac{1}{T\log(n)}}$ and $\alpha\approx\frac{\beta}{\sqrt{T}}\approx\frac{1}{T\sqrt{\log(n)}}$, and get a margin guarantee of approximately $\frac{\alpha}{\sqrt{T}}\approx\frac{1}{T\sqrt{T\log(n)}}$.
A bit more formally, pick (say) $\delta=1/10$ in both Proposition 4.6 and Lemma 4.5, and choose $\beta=\frac{1}{c\sqrt{T\log(n)}}$ for some sufficiently large universal constant $c$, as well as $\alpha=\frac{\beta}{4\sqrt{T}}=\frac{1}{4cT\sqrt{\log(n)}}$. It is easily verified that with this choice, the assumptions in Proposition 4.6, Lemma 4.4 and Lemma 4.5 are all satisfied, and (with a union bound), the resulting matrix $A$ satisfies both
$\max_{l\in[n]}\|A_{l}\|\leq 1~,~~~\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq 0~,\text{~~~and~~~}\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{\alpha}{\sqrt{T+1}}~\geq\frac{1}{c^{\prime}T^{3/2}\sqrt{\log(n)}}$
(for some universal constant $c^{\prime}>0$) with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist.
Appendix B Missing proofs from Section 5
B.1 Proof of Proposition 5.1
@@ Appendix B Missing proofs from Section 5
B.2 Proof of Lemma 5.3
Consider $\mathbf{w}=\frac{1}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which by construction satisfies $\|\mathbf{w}\|_{1}\leq 1$. Since $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthonormal, we have
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{1}\cdot A\mathbf{w}$~ $=~\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}+\mathbf{v}_{T}\mathbf{u}_{T+1}^{\top}\right)\left(\sum_{t=1}^{T+1}\mathbf{u}_{t}\right)$
$=~\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})+\mathbf{v}_{T}~=~\mathbf{v}_{0}~=~\alpha\cdot\mathbf{1}~.$ (14)
Moreover, it holds that
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{1}\leq\sqrt{d}\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{2}=\sqrt{d(T+1)}~,$
and so by Eq. (14), each entry of $A\mathbf{w}$ is at least $\alpha/\sqrt{d(T+1)}$. This holds for the $\mathbf{w}$ we have chosen, so the maximum of $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ over all vectors $\mathbf{w}$ in the unit $\ell_{1}$ ball can only be larger, from which the lemma follows.
B.3 Proof of Lemma 5.4
For any $l\in[n],i\in[d]$ it holds that
$|A_{l,i}|=\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}+v_{T,l}u_{T+1,i}\right|\leq\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}\right|+|v_{T,l}|\cdot|u_{T+1,i}|~.$ (15)
We turn to bound the terms above. For $l\in[n],i\in[d]$, we denote $D^{j}_{l}:=v_{j-1,l}-v_{j,l}$ and $S_{l,i}:=\sum_{j=1}^{T}D^{j}_{l}u_{j,i}$. Note that conditioned on $\mathbf{v}_{0},\dots,\mathbf{v}_{T}$, each $u_{j,i}$ is a sub-Gaussian, mean-zero random variable with sub-Gaussian norm bounded by $O(1/d)$, and therefore $S_{l,i}$ has sub-Gaussian norm at most $O(\frac{1}{d})\cdot\sum_{j=1}^{T}(D_{l}^{j})^{2}$ (cf. Vershynin, 2018 for background on sub-Gaussian random variables). In other words, there exists a universal constant $C_{1}>0$ such that for any $z>0:$
The numerical inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ implies that $\sum_{j=1}^{T}(D^{j}_{l})^{2}\leq 4\sum_{j=0}^{T}v_{j,l}^{2}=4\alpha^{2}+4\sum_{j=1}^{T}v_{j,l}^{2}$, and further note that while $v_{1,l},\dots,v_{T,l}$ are not independent, it does hold that conditionally $v_{t,l}\mid v_{0,l},\dots v_{t-1,l}\sim\mathcal{N}(0,\sigma_{t,l}^{2})$ for some $\sigma_{t,l}\leq\beta$ since projecting a Gaussian cannot increase its variance. Therefore, Lemma C.2 ensures that with probability at least $1-\delta:~\sum_{j=1}^{T}v_{j,l}^{2}\leq 8\beta^{2}(T+\log(1/\delta))$, so via a union bound, we get that with probability at least $1-2\delta$ it holds for all $z>0$ that
Equivalently, choosing $z$ by equating the above to $\delta$, slightly simplifying the expression and union bounding over $l\in[n]$ and $i\in[d]$, we see that for some sufficiently large absolute constant $C>0$ it holds with probability at least $1-3nd\delta$ that
$\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}\right|=|S_{l,i}|\leq C\sqrt{\frac{\log(1/\delta)}{d}}\left[\alpha+\beta(\sqrt{T}+\sqrt{\log(1/\delta)})\right]~,$ (16)
for all $l\in[n],i\in[d]$ simultaneously.
To bound the second term in Eq. (15), by applying Lemma C.1 (with appropriately chosen $z$) it follows that for any $\delta\in(0,1)$, with probability at least $1-\delta:$
$\|\mathbf{v}_{T}\|_{\infty}~\leq~\beta\sqrt{2\log(2n/\delta)}~.$ (17)
Also, applying the same lemma with respect to $\mathbf{u}_{T+1}$, we have that with probability at least $1-\left(\delta+\exp(-d/48)\right):$
$\|\mathbf{u}_{T+1}\|_{\infty}\leq\sqrt{\frac{8\log(2d/\delta)}{d}}~.$ (18)
Overall, under the desired events of Eqs. (16), (17) and (18), we see by Eq. (15) for all $l\in[n],i\in[d]:$
$|A_{l,i}|$ $\leq C\sqrt{\frac{\log(1/\delta)}{d}}[\alpha+\beta(\sqrt{T}+\sqrt{\log(1/\delta)})]+4\beta\sqrt{\frac{\log(2n/\delta)\log(2d/\delta)}{d}}$
$\leq C\frac{\log((n+d)/\delta)}{\sqrt{d}}\left(\alpha+\beta\sqrt{T}+1\right)~,$
the latter holding as long as $C$ is a sufficiently large absolute constant. For any $\delta\in(0,1)$ such that the expression above is at most $1$, which is equivalent to the assumption stated in the lemma, we see that $A\in[-1,1]^{n\times d}$ as required. All this holds with probability at least $1-3nd\delta-2\delta-\exp(-d/48)$ by a union bound, which can be coarsely lower bounded by $1-5nd\delta-\exp(-d/48)$.
B.4 Proof of Theorem 5.2
We construct the matrix $A$ as in Construction 2. Examining the conditions in Proposition 4.6, and Lemma 5.3 and Lemma 5.4, and ignoring log factors momentarily, we see that in order for all of them to be applicable, we must choose $\alpha\lesssim\frac{\beta}{\sqrt{T}}$ (Proposition 4.6), satisfy $\beta\lesssim\sqrt{d/T}$ and $\alpha\lesssim\sqrt{d}$ (Lemma 5.4), yet the game value guarantee is $\approx\frac{\alpha}{\sqrt{Td}}$ (Lemma 5.3). Thus, to make the game value $\frac{\alpha}{\sqrt{Td}}$ as large as possible, yet satisfying all of the constraints, we should choose $\beta\approx\sqrt{d/T}$ and $\alpha\approx\beta/\sqrt{T}\approx\sqrt{d}/T$, and get a margin guarantee of approximately $\frac{\alpha}{\sqrt{Td}}\approx\frac{1}{T^{3/2}}$.
A bit more formally, pick (say) $\delta=1/(50nd)$ in both Proposition 4.6, Lemma 5.3 and Lemma 5.4, and choose
$\beta=\frac{\sqrt{d}}{3C\sqrt{T}\log(50nd(n+d))}$
for the sufficiently large universal constant $C$ given by Lemma 5.4, as well as
$\alpha=\frac{\beta}{4\sqrt{T}}=\frac{\sqrt{d}}{12CT\log(50nd(n+d))}~.$
It is easily verified that with this choice, as well as the theorem assumptions, the conditions in Proposition 4.6, Lemma 5.3 and Lemma 5.4 are all satisfied, and (with a union bound), the resulting matrix $A$ satisfies both $A\in[-1,+1]^{n\times d}$ and $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq~0$, as well as (for some constant $c^{\prime}>0$)
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}\geq\frac{\alpha}{\sqrt{d(T+1)}}=\frac{\beta}{4\sqrt{T}}=\frac{1}{12C\log(50nd(n+d))}\cdot\frac{1}{T^{3/2}}$
all with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist. Coarsely upper bounding $\log(50nd(n+d))\leq\log(100n^{2}d^{2})=2\log(10nd)$, and plugging in the displayed equation above, the theorem follows.
Appendix C Auxiliary lemmas
Lemma C.1.
@@ Appendix C Auxiliary lemmas
As for the second inequality, note that since $M$ has at most $q/2$ orthonormal columns, then $I-MM^{\top}$ is a projection matrix to a subspace of $\mathbb{R}^{q}$ of dimensionality at least $q/2$. Thus, $\|(I-MM^{\top})\boldsymbol{\xi}\|$ has the same distribution as the norm of a standard Gaussian random variable on $\mathbb{R}^{s}$ where $s\geq q/2$. Using a standard tail lower bound for such Gaussian norms (see for example [Shalev-Shwartz and Ben-David, 2014, Lemma B.12]), it follows that

### v4

@@ 7 Proofs
This bound is true for any fixed unit $\mathbf{r}$. In order to extend the bound simultaneously for all unit $\mathbf{r}$, we will use a standard $\epsilon$-net argument: Fix some $\epsilon>0$, and let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net of the unit Euclidean ball in $\mathbb{R}^{T}$, of size $\leq\left(\frac{2}{\epsilon}+1\right)^{T}$ (namely, a collection of unit vectors so that each $\mathbf{r}$ in the unit ball is $\epsilon$-close in Euclidean distance to one of the vectors). The existence of an $\epsilon$-net of this size is shown, for example, in (Vershynin, 2012, Lemma 5.2). Using a union bound and the displayed equation above, it follows that with probability at least $1-\left(\frac{2}{\epsilon}+1\right)^{T}\exp\left(-\frac{n}{30}\right)$,
@@
Combining Lemma 7.1 and Lemma 7.2 with a union bound, and applying them to Eq. (11), we get the following: With probability at least $1-\delta-\exp\left(T\log(2n)-\frac{n}{32}\right)$ over $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{T}$, we have that for any $\mathbf{w}$, there is some subset $\mathcal{L}\subseteq[n]$ of size $|\mathcal{L}|\geq\frac{n}{20}$ for which the following inequalities hold:
@@
where $(1)$ is by Lemma 7.1, $(2)$ is by the fact that a minimum can be upper bounded by an average, $(3)$ is by Lemma 7.2 and the fact that $|\mathcal{L}|\geq\frac{n}{20}$, $(4)$ is by the assumption in the proposition statement, and $(5)$ is by definition of the vector $\mathbf{r}$ and scalar $r_{0}$ in Eq. (10).
@@
7.6 Proof of Proposition 6.1
@@
7.7 Auxiliary lemmas
Lemma 7.4.
@@ 7.7 Auxiliary lemmas
As for the second inequality, note that since $M$ has at most $q/2$ orthonormal columns, then $I-MM^{\top}$ is a projection matrix to a subspace of $\mathbb{R}^{q}$ of dimensionality at least $q/2$. Thus, $\|(I-MM^{\top})\boldsymbol{\xi}\|$ has the same distribution as the norm of a standard Gaussian random variable on $\mathbb{R}^{s}$ where $s\geq q/2$. Using a standard tail lower bound for such Gaussian norms (see for example (Shalev-Shwartz and Ben-David, 2014, Lemma B.12)), it follows that
@@
8 Discussion
In this paper, we studied the oracle complexity of solving matrix games where (at least) one of the constraint sets is the simplex. This well-studied problem class encapsulates canonical tasks, including finding a linear separator and computing a Nash equilibrium in bilinear zero-sum games, as well as some others discussed throughout the paper. By identifying distinct oracle models corresponding to either one-sided or two-sided multiplication by the matrix $A$, we were able to shed light on previous algorithmic approaches for this task, and to provide several new lower bounds.
The most notable implications of our work are that alongside the recent upper bounds of Karmarkar et al. (2026), our work resolves the complexity of these two canonical tasks on the order of $\tilde{\Theta}(\epsilon^{-2/3})$ (up to log factors), in the natural oracle model in which the interaction with the matrix is granted through multiplications from both sides. Moreover, our lower bound in the one-sided multiplication model points out that any acceleration beyond the classic $O(\epsilon^{-2})$ rate provably requires such two-sided operations.
We note that extending our lower bounds to randomized algorithms remains an open problem. Since our constructions required simulation of the algorithm’s responses, it is not immediately clear how this can be achieved. We further note that previous extensions of oracle complexity lower bounds to randomized algorithms (Nemirovski and Yudin, 1983; Carmon et al., 2020a; Arjevani et al., 2023) crucially relied on non-linear modifications of the “hard” target function, which is not possible when we restrict ourselves to bilinear functions.
Acknowledgments